ALGEBRA 1
Write in standard form: y+5=4/5(x+20)

Answer by mananth(16946)

y+5=4/5(x+20)
multiply by 5
5(y+5)=4(x+20)
5y+25=4x+80
-4x+5y=80-25
-4x+5y=55
...